The heating element of a water heater in an apartment building has a maximum power output of 27 kW. Four residents of the building take showers at the same time, and each receives heated water at a volume flow rate of 12x10-5 m3/s. If the water going into the heater has a temperature of 20°C, what is the maximum possible temperature of the hot water that each showering resident receives?
Answer #1

Volume of Water passing through the heater in 1 sec = 1.2 x 10^-4 m^3
=>mass of water passing through the heater in 1 sec = 1.2 x 10^-4 x 10^3 = 0.12 kg
=>By Q = ms∆t
=>Q = 0.12 x 4187 x ∆t
=>27000 = 0.12 x 4187 x ∆t
=>∆t = 53.74 *C
=>t(final) = 20 + 53.74 = 73.74*C